OpenStudy (anonymous):

[7.04] Choose the correct simplification of f to the 9th power times h to the 23rd power all over f to the 3rd power times h to the 17th power. Select one: a. f12h6 b. 1 over f to the 12th power times h to the 6th power c. f6h6 d. 1 over f to the 6th power times h to the 6th power

OpenStudy (johnweldon1993):

\[\huge \frac{f^9 \times h^{23}}{f^3 \times h^{17}} = f^{9 - 3} \times h^{23 - 17} = \space ?\]

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

chewbacca

OpenStudy (anonymous):

@johnweldon1993